# Representation Learning for Distributional Perturbation Extrapolation

### v1

@@ Appendix A Proofs
By Thm. 4.1, we have $\mathbf{h}(\mathbf{z})=\mathbf{A}\mathbf{z}+\mathbf{b}$ for
$\mathbf{A}$ $=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}},$ (A.25)
$\mathbf{b}$ $=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}.$ (A.26)
Together with $\mathbf{Z}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$, this lets us compute the distribution of the LHS of (A.24) as:
$\mathbf{h}(\mathbf{Z}+\mathbf{W}\mathbf{l}_{\mathrm{test}})=\mathbf{A}(\mathbf{Z}+\mathbf{W}\mathbf{l}_{\mathrm{test}})+\mathbf{b}\sim\mathcal{N}\left(\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b},\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}\right).$ (A.27)
Similarly, since $\widetilde{\mathbf{Z}}\sim\mathcal{N}\left(\widetilde{\bm{\mu}},\widetilde{\bm{\Sigma}}\right)$, the distribution of the RHS of (A.24) is given by
$\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}}\sim\mathcal{N}\left(\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}},\widetilde{\bm{\Sigma}}\right).$ (A.28)
From (A.25), it follows directly that
$\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}=\widetilde{\bm{\Sigma}}.$ (A.29)
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}}.$ (A.30)
Starting from the LHS of (A.30), by substituting the expression for $\mathbf{b}$ from (A.26) we obtain:
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}$ $=\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}$ (A.31)
$=\widetilde{\bm{\mu}}+\mathbf{A}\mathbf{W}(\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0})+\widetilde{\mathbf{W}}\mathbf{l}_{0}$ (A.32)
Next, by (4.6), i.e., the assumption that $(\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0})$ lies in the span of $\{\mathbf{l}_{e}-\mathbf{l}_{0}\}_{e\in[M]}$, there exists $\bm{\alpha}\in\mathbb{R}^{M}$ such that
$\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0}=\sum_{e\in[M]}\alpha_{e}(\mathbf{l}_{e}-\mathbf{l}_{0})=\mathbf{L}\bm{\alpha}$ (A.33)
where, as before, $\mathbf{L}\in\mathbb{R}^{K\times M}$ is the matrix with columns $(\mathbf{l}_{e}-\mathbf{l}_{0})$. Substituting (A.33) into (A.32) yields
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}=\widetilde{\bm{\mu}}+\mathbf{A}\mathbf{W}\mathbf{L}\bm{\alpha}+\widetilde{\mathbf{W}}\mathbf{l}_{0}.$ (A.34)
Finally, it follows from Thm. 4.1 that
$\mathbf{A}\mathbf{W}\mathbf{L}=\widetilde{\mathbf{W}}\mathbf{L}.$ (A.35)
which upon substituion into (A.34) yields the desired equality from (A.30)
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}$ $=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{L}\bm{\alpha}+\widetilde{\mathbf{W}}\mathbf{l}_{0}$ (A.36)
$=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}(\mathbf{L}\bm{\alpha}+\mathbf{l}_{0})$ (A.37)
$=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}}.$ (A.38)
This completes the proof. ∎
Appendix B Additional Background Material and Related Work
@@
Since the problem of interest (§ 2) involves making distributional predictions for new perturbation conditions, we review some basics of probabilistic forecasting (§ B.1) and measuring the similarity between two distributions (§ B.2), in our case typically between an empirical distribution and its predicted counterpart. We then turn to representation learning with encoder-decoder architectures (§ B.3), in particular a recent approach that also targets distributional reconstruction. Finally, we cover some prior efforts on perturbation modelling and extrapolation (§ B.4), which we draw inspiration from.
B.1 Probabilistic Forecasting and Scoring Rules
@@
CRPS.
For continuous scalar random variables (i.e., $\Omega=\mathbb{R}$), a popular scoring rule is the continuous ranked probability score (CRPS; Matheson & Winkler, 1976). When $\mathcal{P}$ is the space of Borel probability measures on $\mathbb{R}$ with finite first moment, it is strictly proper and given by: [^4] The original definition by Matheson & Winkler (1976) is $\operatorname{CRPS}(F,x)=-\int_{-\infty}^{\infty}(F(y)-1\{y\geq x\})^{2}\mathrm{d}y$ where $F$ is the CDF of $\mathbb{P}$, but the simpler form in (B.1) has been shown to be equivalent (Baringhaus & Franz, 2004, Lemma 2.2).
[…]
@@
Energy Score.
Gneiting & Raftery (2007) propose the energy score as a multi-variate generalization of the CRPS for vector-valued $\mathbf{x}\in\Omega=\mathbb{R}^{{d_{X}}}$. For $\beta\in(0,2)$, it is defined by
[…]
@@
B.2 Assessing Distributional Similarity
Energy Distance.
The expected energy score $\operatorname{ES}_{\beta}(\mathbb{P},\mathbb{Q})=\mathbb{E}_{\mathbf{Y}\sim\mathbb{Q}}[\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{Y})]$ consitutes a measure of similarity between $\mathbb{P}$ and $\mathbb{Q}$ and is closely linked to the energy distance (Székely & Rizzo, 2013):
$\operatorname{ED}_{\beta}(\mathbb{P},\mathbb{Q})$ $=2\mathbb{E}_{\mathbf{X}\sim\mathbb{P},\mathbf{Y}\sim\mathbb{Q}}\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{X}^{\prime}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{Y},\mathbf{Y}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}}\left\lVert\mathbf{Y}-\mathbf{Y}^{\prime}\right\rVert_{2}^{\beta}$ (B.3)
$=2\left(\operatorname{ES}_{\beta}(\mathbb{Q},\mathbb{Q})-\operatorname{ES}_{\beta}(\mathbb{P},\mathbb{Q})\right)\geq 0$ (B.4)
@@ B.2 Assessing Distributional Similarity
Maximum Mean Discrepancy (MMD).
Another well-known distance between probability measures that is rooted in kernel methods (Schölkopf & Smola, 2002) is the maximum mean discrepancy (MMD; Gretton et al., 2012), which for a positive definite kernel $k:\Omega\times\Omega\to\mathbb{R}$ is given by
$\operatorname{MMD}^{2}_{k}(\mathbb{P},\mathbb{Q})=\mathbb{E}_{X,X^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{P}}[k(X,X^{\prime})]-2\mathbb{E}_{X\sim\mathbb{P},Y\sim\mathbb{Q}}[k(X,Y)]+\mathbb{E}_{Y,Y^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}}[k(Y,Y^{\prime})]\,.$ (B.5)
Energy Distance as a Special Case of MMD.
As shown by Sejdinovic et al. (2013), if $k$ in (B.5) is chosen to be the positive-definite distance kernel555induced by the negative-definite semi-metric $\rho(\mathbf{X},\mathbf{Y})=\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}$ on $\Omega=\mathbb{R}^{{d_{X}}}$ (centered at the origin),
$k_{\beta}(\mathbf{X},\mathbf{Y})=\frac{1}{2}\left(\left\lVert\mathbf{X}\right\rVert_{2}^{\beta}+\left\lVert\mathbf{Y}\right\rVert_{2}^{\beta}-\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}\right)$ (B.6)
@@
$\operatorname{ED}_{\beta}(\mathbb{P},\mathbb{Q})=2\operatorname{MMD}^{2}_{k_{\beta}}(\mathbb{P},\mathbb{Q})\,.$ (B.7)
B.3 Representation Learning
@@
Autoencoder (AE).
An autoencoder (AE; Hinton & Salakhutdinov, 2006; Rumelhart et al., 1986) is a pair of functions $(\mathbf{g},\mathbf{f})$, consisting of an encoder $\mathbf{g}:\mathbb{R}^{{d_{X}}}\to\mathbb{R}^{{d_{Z}}}$ mapping observations $\mathbf{X}\sim\mathbb{P}$ to their representation $\mathbf{Z}:=\mathbf{g}(\mathbf{X})$, and a decoder $\mathbf{f}:\mathbb{R}^{{d_{Z}}}\to\mathbb{R}^{{d_{X}}}$ mapping representations $\mathbf{Z}$ to their reconstructions in observation space, $\widehat{\mathbf{X}}:=\mathbf{f}(\mathbf{Z})=\mathbf{f}(\mathbf{g}(\mathbf{X}))$. Typically, ${d_{Z}}<{d_{X}}$, such that there is a bottleneck and perfect reconstruction is not feasible. Autoencoders are usually trained with a (point-wise, mean) reconstruction objective, i.e., the aim is to minimise the mean squared error
$\mathcal{L}_{\textsc{ae}}(\mathbf{g},\mathbf{f};\mathbb{P}):=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\widehat{\mathbf{X}}\right\rVert_{2}^{2}=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{f}(\mathbf{g}(\mathbf{X}))\right\rVert_{2}^{2}\,,$ (B.8)
@@
$\mathbf{f}^{*}_{\textsc{ae}}(\mathbf{z};\mathbf{g},\mathbb{P})=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}[\mathbf{X}|\mathbf{g}(\mathbf{X})=\mathbf{z}]\,.$ (B.9)
Distributional Principal Autoencoder (DPA).
Since the objective of a standard AE is mean reconstruction, the distribution of reconstructions $\widehat{\mathbf{X}}$ is typically not the same as the distribution of $\mathbf{X}$ (unless the encoder is invertible, i.e, the compression is lossless and perfect reconstruction is feasible, which is usually not the case in practice). To address this, Shen & Meinshausen (2024a) proposed the distributional principal autoencoder (DPA) which targets distributional (rather than mean) reconstruction. A DPA also consists of an encoder-decoder pair $(\mathbf{g},\mathbf{f})$. However, unlike in a standard AE, the DPA decoder $\mathbf{f}:\mathbb{R}^{{d_{Z}}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{{d_{X}}}$ is stochastic and receives an additional noise term ${\bm{\varepsilon}}$ as input, which is sampled from a fixed distribution $\mathbb{Q}_{{\bm{\varepsilon}}}$ such as a standard isotropic Gaussian. The DPA loss function is constructed such that for a fixed encoder $\mathbf{g}$, the optimal DPA decoder $\mathbf{f}^{*}_{\textsc{dpa}}$ maps a given latent embedding $\mathbf{z}$, to the distribution of $\mathbf{X}$, given $\mathbf{g}(\mathbf{X})=\mathbf{z}$, i.e.,
[…]
$\mathcal{L}_{\textsc{dpa}}(\mathbf{g},\mathbf{f};\mathbb{P})$ $=\mathbb{E}_{\mathbf{X}\sim\mathbb{P},{\bm{\varepsilon}}\sim\mathbb{Q}_{{\bm{\varepsilon}}}}\left\lVert\mathbf{X}-\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}})\right\rVert_{2}^{\beta}-\frac{1}{2}\mathbb{E}_{\mathbf{X}\sim\mathbb{P},{\bm{\varepsilon}},{\bm{\varepsilon}}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}_{{\bm{\varepsilon}}}}\left\lVert\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}})-\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}}^{\prime})\right\rVert_{2}^{\beta}$ (B.11)
@@
B.4 Perturbation Modelling
Compositional Perturbation Autoencoder (CPA).
Lotfollahi et al. (2023) propose the compositional perturbation autoencoder (CPA) as a model for compositional extrapolation of perturbation data. Specifically, they assume the following model:
$\mathbf{z}^{\mathrm{pert}}$ $=\mathbf{z}^{\mathrm{base}}+\mathbf{W}^{\mathrm{pert}}\begin{pmatrix}h_{1}(l_{1})\\ \dots\\ h_{K}(l_{K})\end{pmatrix}+\sum_{j=1}^{J}\mathbf{W}_{j}^{\text{cov}}\mathbf{c}_{j}\,,$ (B.12)
where $\mathbf{z}^{\mathrm{base}}\sim\mathbb{P}_{\mathbf{Z}}$ denotes an unperturbed basal state; the matrix $\mathbf{W}^{\mathrm{pert}}\in\mathbb{R}^{{d_{Z}}\times K}$ encodes the additive effect of each elementary perturbation; $\{h_{k}:\mathbb{R}\to\mathbb{R}\}_{k\in[K]}$ are unknown, possibly nonlinear dose-response curves; $\{\mathbf{c}_{j}\in\mathbb{R}^{K_{j}}\}_{j\in[J]}$ are observed one-hot vectors capturing $J$ additional discrete covariates, such as cell-types or species; and the matrices $\{\mathbf{W}_{j}^{\text{cov}}\in\mathbb{R}^{{d_{Z}}\times K_{j}}\}_{j\in[J]}$ encode additive covariate-specific effects. Further, the basal state $\mathbf{z}^{\mathrm{base}}$ is assumed to be independent of the perturbation labels $\mathbf{l}$ and covariates $\mathbf{C}=(\mathbf{c}_{1},\dots,\mathbf{c}_{J})$. Observations $\mathbf{x}$ are then drawn from a Gaussian whose mean and variance are determined by the perturbed latent state $\mathbf{z}^{\mathrm{pert}}$,
$\mathbf{x}\sim\mathcal{N}\left(\bm{\mu}\left(\mathbf{z}^{\mathrm{pert}}\right),\sigma^{2}\left(\mathbf{z}^{\mathrm{pert}}\right)\mathbf{I}\right)\,.$ (B.13)

### v2

@@ Appendix B PDAE training: Additional losses and complete algorithm
[…]
2:for $i=1,\dots,B$ do
3:  $\widehat{\mathbf{z}}^{\mathrm{pert}}_{i}\leftarrow\widehat{\mathbf{g}}_{\bm{\phi}}(\mathbf{x}_{i})$ # encode
4:  ${\bm{\varepsilon}}_{i},{\bm{\varepsilon}}_{i}^{\prime}\sim\mathcal{N}(\bm{0},\mathbf{I})$ # noise for reconstructions
5:  $\widehat{\mathbf{x}}_{i},\widehat{\mathbf{x}}_{i}^{\prime}\leftarrow\widehat{\mathbf{f}}_{\bm{\theta}}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{i},{\bm{\varepsilon}}_{i}\right),\widehat{\mathbf{f}}_{\bm{\theta}}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{i},{\bm{\varepsilon}}_{i}^{\prime}\right)$ # noisy reconstructions
6:  $\mathcal{L}_{\text{cond-rec}}({\bm{\phi}},{\bm{\theta}})\mathrel{+}=\frac{1}{2B}\left(\left\lVert\mathbf{x}_{i}-\widehat{\mathbf{x}}_{i}\right\rVert+\left\lVert\mathbf{x}_{i}-\widehat{\mathbf{x}}_{i}^{\prime}\right\rVert-\left\lVert\widehat{\mathbf{x}}_{i}-\widehat{\mathbf{x}}_{i}^{\prime}\right\rVert\right)$ # conditional energy score
7:  $\widehat{\mathbf{z}}^{\mathrm{base}}_{i}\leftarrow\widehat{\mathbf{z}}_{i}^{\mathrm{pert}}-\widehat{\mathbf{W}}\mathbf{a}_{i}$ # estimated basal states
8:  for $j\in[B]\setminus\{i\}$ do
9:   $\widehat{\mathbf{z}}^{\mathrm{pert}}_{j\to i}\leftarrow\mathbf{z}_{j}^{\mathrm{base}}+\widehat{\mathbf{W}}\mathbf{a}_{i}$ # synthetic perturbed latents
10:   ${\bm{\varepsilon}}_{ji}\sim\mathcal{N}(\bm{0},\mathbf{I})$ # noise for synthetic observations
11:   $\widehat{\mathbf{x}}_{j\to i}\leftarrow\widehat{\mathbf{f}}_{\theta}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{j\to i},{\bm{\varepsilon}}_{ji}\right)$ # synthetic observations
12:  end for
13:  $\mathcal{L}_{\mathrm{pert}}\left({\bm{\phi}},{\bm{\theta}},\widehat{\mathbf{W}}\right)\mathrel{+}=\frac{1}{B(B-1)}\sum_{j\in[B]\setminus\{i\}}\left[\left\lVert\mathbf{x}_{i}-\widehat{\mathbf{x}}_{j\to i}\right\rVert-\frac{1}{B-2}\sum_{i\neq j^{\prime}<j}\left\lVert\widehat{\mathbf{x}}_{j\to i}-\widehat{\mathbf{x}}_{j^{\prime}\to i}\right\rVert\right]$
14:  $\bm{\xi}_{i}\sim\mathcal{N}\left(\bm{0},\mathbf{I}\right)$ # noise for prior loss
15:end for
[…]
MMD-based goodness of fit assessment. To assess goodness of fit (e.g., for early stopping or to assess training convergence), we propose to assess the pairwise distributional fits between simulated and true perturbation distributions using MMD. That is, to consider $\operatorname{MMD}(\mathbb{P}_{e\to h},\mathbb{P}_{e})$ for all $e,h\in[M]_{0}$.
Code. We will publicly release PDAE code upon publication.
Appendix C Additional background material and related work
@@
Since the problem of interest (§˜2) involves making distributional predictions for new perturbation conditions, we review some basics of probabilistic forecasting (§˜C.1) and measuring the similarity between two distributions (§˜C.2), in our case typically between an empirical distribution and its predicted counterpart. We then turn to representation learning with encoder-decoder architectures (§˜C.3), in particular a recent approach that also targets distributional reconstruction. Finally, we cover some prior efforts on perturbation modeling and extrapolation (§˜C.4), which we draw inspiration from.
C.1 Probabilistic forecasting and scoring rules
@@
CRPS. For continuous scalar random variables (i.e., $\Omega=\mathbb{R}$), a popular scoring rule is the continuous ranked probability score (CRPS; Matheson & Winkler, 1976). When $\mathcal{P}$ is the space of Borel probability measures on $\mathbb{R}$ with finite first moment, it is strictly proper and given by: [^4] The original definition by Matheson & Winkler (1976) is $\operatorname{CRPS}(F,x)=-\int_{-\infty}^{\infty}(F(y)-1\{y\geq x\})^{2}\mathrm{d}y$ where $F$ is the CDF of $\mathbb{P}$, but the simpler form in (C.1) has been shown to be equivalent (Baringhaus & Franz, 2004, Lemma 2.2).
$\operatorname{CRPS}(\mathbb{P},x)=\frac{1}{2}\mathbb{E}_{X,X^{\prime}\,\overset{\text{i.i.d.}}{\sim}\,\mathbb{P}}\left|X-X^{\prime}\right|-\mathbb{E}_{X\sim\mathbb{P}}\left|X-x\right|\,.$ (C.1)
@@
Energy score. Gneiting & Raftery (2007) propose the energy score as a multi-variate generalization of the CRPS for vector-valued $\mathbf{x}\in\Omega=\mathbb{R}^{{d_{X}}}$. For $\beta\in(0,2)$, it is defined by
$\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{x})=\frac{1}{2}\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\,\overset{\text{i.i.d.}}{\sim}\,\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{X}^{\prime}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{x}\right\rVert_{2}^{\beta}\,$, (C.2)
@@
C.2 Assessing distributional similarity
Energy distance. The expected energy score $\operatorname{ES}_{\beta}(\mathbb{P},\mathbb{Q})=\mathbb{E}_{\mathbf{Y}\sim\mathbb{Q}}[\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{Y})]$ consitutes a measure of similarity between $\mathbb{P}$ and $\mathbb{Q}$ and is closely linked to the energy distance (Székely & Rizzo, 2013):
$\operatorname{ED}_{\beta}(\mathbb{P},\mathbb{Q})$ $=2\mathbb{E}_{\mathbf{X}\sim\mathbb{P},\mathbf{Y}\sim\mathbb{Q}}\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{X}^{\prime}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{Y},\mathbf{Y}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}}\left\lVert\mathbf{Y}-\mathbf{Y}^{\prime}\right\rVert_{2}^{\beta}$ (C.3)
$=2\left(\operatorname{ES}_{\beta}(\mathbb{Q},\mathbb{Q})-\operatorname{ES}_{\beta}(\mathbb{P},\mathbb{Q})\right)\geq 0$ (C.4)
@@ C.2 Assessing distributional similarity
Maximum mean discrepancy (MMD). Another well-known distance between probability measures that is rooted in kernel methods (Schölkopf & Smola, 2002) is the maximum mean discrepancy (MMD; Gretton et al., 2012), which for a positive definite kernel $k:\Omega\times\Omega\to\mathbb{R}$ is given by
$\operatorname{MMD}^{2}_{k}(\mathbb{P},\mathbb{Q})=\mathbb{E}_{X,X^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{P}}[k(X,X^{\prime})]-2\mathbb{E}_{X\sim\mathbb{P},Y\sim\mathbb{Q}}[k(X,Y)]+\mathbb{E}_{Y,Y^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}}[k(Y,Y^{\prime})]\,.$ (C.5)
Energy distance as a special case of MMD. As shown by Sejdinovic et al. (2013), if $k$ in (C.5) is chosen to be the positive-definite distance kernel555induced by the negative-definite semi-metric $\rho(\mathbf{X},\mathbf{Y})=\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}$ on $\Omega=\mathbb{R}^{{d_{X}}}$ (centered at the origin),
$k_{\beta}(\mathbf{X},\mathbf{Y})=\frac{1}{2}\left(\left\lVert\mathbf{X}\right\rVert_{2}^{\beta}+\left\lVert\mathbf{Y}\right\rVert_{2}^{\beta}-\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}\right)$ (C.6)
@@
$\operatorname{ED}_{\beta}(\mathbb{P},\mathbb{Q})=2\operatorname{MMD}^{2}_{k_{\beta}}(\mathbb{P},\mathbb{Q})\,.$ (C.7)
C.3 Representation learning
@@
Autoencoder (AE). An autoencoder (AE; Hinton & Salakhutdinov, 2006; Rumelhart et al., 1986) is a pair of functions $(\mathbf{g},\mathbf{f})$, consisting of an encoder $\mathbf{g}:\mathbb{R}^{{d_{X}}}\to\mathbb{R}^{{d_{Z}}}$ mapping observations $\mathbf{X}\sim\mathbb{P}$ to their representation $\mathbf{Z}:=\mathbf{g}(\mathbf{X})$, and a decoder $\mathbf{f}:\mathbb{R}^{{d_{Z}}}\to\mathbb{R}^{{d_{X}}}$ mapping representations $\mathbf{Z}$ to their reconstructions in observation space, $\widehat{\mathbf{X}}:=\mathbf{f}(\mathbf{Z})=\mathbf{f}(\mathbf{g}(\mathbf{X}))$. Typically, ${d_{Z}}<{d_{X}}$, such that there is a bottleneck and perfect reconstruction is not feasible. Autoencoders are usually trained with a (point-wise, mean) reconstruction objective, i.e., the aim is to minimise the mean squared error
$\mathcal{L}_{\textsc{ae}}(\mathbf{g},\mathbf{f};\mathbb{P}):=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\widehat{\mathbf{X}}\right\rVert_{2}^{2}=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{f}(\mathbf{g}(\mathbf{X}))\right\rVert_{2}^{2}\,$, (C.8)
@@
$\mathbf{f}^{*}_{\textsc{ae}}(\mathbf{z};\mathbf{g},\mathbb{P})=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}[\mathbf{X}|\mathbf{g}(\mathbf{X})=\mathbf{z}]\,.$ (C.9)
Distributional principal autoencoder (DPA). Since the objective of a standard AE is mean reconstruction, the distribution of reconstructions $\widehat{\mathbf{X}}$ is typically not the same as the distribution of $\mathbf{X}$ (unless the encoder is invertible, i.e, the compression is lossless and perfect reconstruction is feasible, which is usually not the case in practice). To address this, Shen & Meinshausen (2024a) proposed the distributional principal autoencoder (DPA) which targets distributional (rather than mean) reconstruction. A DPA also consists of an encoder-decoder pair $(\mathbf{g},\mathbf{f})$. However, unlike in a standard AE, the DPA decoder $\mathbf{f}:\mathbb{R}^{{d_{Z}}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{{d_{X}}}$ is stochastic in that it receives an additional noise term $\bm{\varepsilon}$} as input, which is distributed according to a fixed distribution $\mathbb{Q}_{{\bm{\varepsilon}}}$ such as a standard isotropic Gaussian. The DPA loss function is constructed such that for a fixed encoder $\mathbf{g}$, the optimal DPA decoder $\mathbf{f}^{*}_{\textsc{dpa}}$ maps a given latent embedding $\mathbf{z}$ to the distribution of $\mathbf{X}$, given $\mathbf{g}(\mathbf{X})=\mathbf{z}$, i.e.,
$\mathbf{f}^{*}_{\textsc{dpa}}(\mathbf{z},{\bm{\varepsilon}};\mathbf{g})\overset{d}{=}\left(\mathbf{X}|\mathbf{g}(\mathbf{X})=\mathbf{z}\right)\,$, (C.10)
where $\overset{d}{=}$ denotes equality in distribution. This means that the decoder evaluated at $\mathbf{z}$ should match the distribution of all realizations of $\mathbf{X}$ that are mapped by the encoder to $\mathbf{z}$. At the same time, the DPA encoder aims to minimise the variability in the distributions in (C.10) by encoding the first $d_{Z}$} “principal” components. As shown by Shen & Meinshausen (2024a), both of these goals are achieved by the following DPA objective,
$\mathcal{L}_{\textsc{dpa}}(\mathbf{g},\mathbf{f};\mathbb{P})$ $=\mathbb{E}_{\mathbf{X}\sim\mathbb{P},{\bm{\varepsilon}}\sim\mathbb{Q}_{{\bm{\varepsilon}}}}\left\lVert\mathbf{X}-\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}})\right\rVert_{2}^{\beta}-\frac{1}{2}\mathbb{E}_{\mathbf{X}\sim\mathbb{P},{\bm{\varepsilon}},{\bm{\varepsilon}}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}_{{\bm{\varepsilon}}}}\left\lVert\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}})-\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}}^{\prime})\right\rVert_{2}^{\beta}$ (C.11)
@@
C.4 Perturbation modeling
Compositional perturbation autoencoder (CPA). Lotfollahi et al. (2023) propose the compositional perturbation autoencoder (CPA) as a model for compositional extrapolation of perturbation data. Specifically, they assume the following model:
$\mathbf{z}^{\mathrm{pert}}$ $=\mathbf{z}^{\mathrm{base}}+\mathbf{W}^{\mathrm{pert}}\begin{pmatrix}h_{1}(a_{1})\\ \dots\\ h_{K}(a_{K})\end{pmatrix}+\sum_{j=1}^{J}\mathbf{W}_{j}^{\text{cov}}\mathbf{c}_{j}\,$, (C.12)
where $\mathbf{z}^{\mathrm{base}}\sim\mathbb{P}_{\mathbf{Z}}$ denotes an unperturbed basal state; the matrix $\mathbf{W}^{\mathrm{pert}}\in\mathbb{R}^{{d_{Z}}\times K}$ encodes the additive effect of each elementary perturbation; $\{h_{k}:\mathbb{R}\to\mathbb{R}\}_{k\in[K]}$ are unknown, possibly nonlinear dose-response curves; $\{\mathbf{c}_{j}\in\mathbb{R}^{K_{j}}\}_{j\in[J]}$ are observed one-hot vectors capturing $J$ additional discrete covariates, such as cell-types or species; and the matrices $\{\mathbf{W}_{j}^{\text{cov}}\in\mathbb{R}^{{d_{Z}}\times K_{j}}\}_{j\in[J]}$ encode additive covariate-specific effects. Further, the basal state $\mathbf{z}^{\mathrm{base}}$ is assumed to be independent of the perturbation labels $\mathbf{a}$ and covariates $\mathbf{C}=(\mathbf{c}_{1},\dots,\mathbf{c}_{J})$. Observations $\mathbf{x}$ are then drawn from a Gaussian whose mean and variance are determined by the perturbed latent state $\mathbf{z}^{\mathrm{pert}}$,
$\mathbf{x}\sim\mathcal{N}\left(\bm{\mu}\left(\mathbf{z}^{\mathrm{pert}}\right),\sigma^{2}\left(\mathbf{z}^{\mathrm{pert}}\right)\mathbf{I}\right)\,.$ (C.13)